OpenStudy (anonymous):

what is the exponent form of sqrt (x^3) ?

OpenStudy (jamesj):

\[ \sqrt{y} = y^{1/2} \] hence \[ \sqrt{x^3} = (x^3)^{1/2} = x^{3/2} \]

OpenStudy (anonymous):

yes that's what i got. but if i substitute x=8, the answer for using sqrt (x^3) and x^3/2 is different

OpenStudy (anonymous):

cuz i am suppose to find the slope of the tangent at x=8, which the answer is 3 sqrt 2, but i couldnt get the answer using power rule

OpenStudy (jamesj):

No so. \( 8^3 = 512 \) and \( \sqrt{512} = 16 \sqrt{2} \), while ...

OpenStudy (jamesj):

\( 8^{3/2} = (2^3)^{3/2} = 2^{9/2} = 2^4 \sqrt{2} = 16 \sqrt{2} \)

OpenStudy (jamesj):

Got it?

OpenStudy (anonymous):

umm, but if i do the power rule y' = 3/2 (8) ^1/2 i got some decimal answers

OpenStudy (jamesj):

Ok, now you have \[ y' = \frac{3}{2} \sqrt{8} = \frac{3}{2} 2 \sqrt{2} = 3\sqrt{2} \]

OpenStudy (anonymous):

ohh ok i get it now. so i just cant plug everything in the calculater

OpenStudy (jamesj):

If you started with \( y = \sqrt{x^3} \) and evaluate this with the chain rule, then \[ y' = 3x^2 \cdot \frac{1}{2} \frac{1}{\sqrt{x^3}} \] Work this through, and you'll find it equal to what's above.

OpenStudy (anonymous):

oh ok i will try it! thanks !